OpenStudy (anonymous):

joe has 16 more baseball cards then football cards. he also noticed that of the total he has three times as many baseball cards. how many baseball cards does he have?

OpenStudy (aaronq):

let b = number of baseball cards, f = number of football cards 16 more b cards then f cards: f= b +16 three times as many baseball cards: f=3b make them equal to each other f = 3b = b +16 = f 3b = b +16 -> 3b-b=16 -> 2b=16 -> b=16/2=8 He has 8 baseball cards.